OpenStudy (anonymous):

Horatio has a pack of candies containing 4 strawberry, 2 lemon, 1 lime, and 3 orange flavored chews, in random order. If he eats each candy immediately after taking it out of the pack, which multiplication expression can be used to determine the probability that the first two candies he eats are lemon?

OpenStudy (anonymous):

\[\frac{2}{10}\times \frac{1}{9}=\frac{2}{90}=\frac{1}{45}\]

OpenStudy (anonymous):

how are u soo good at math?

OpenStudy (anonymous):

Always remember that maths is about taking things apart and seeing how they work, not formulas and equations and you'll be good at it.

OpenStudy (anonymous):

Cause I've always wanted to become a fighter pilot and don't know the necessary math to do so

OpenStudy (anonymous):

I didn't know you needed any maths to be a fighter pilot.

OpenStudy (anonymous):

yes, my cousin is a b1 WSO and he told me to b a pilot need to know alot of math

OpenStudy (anonymous):

Oh, really? Do you need a degree?

OpenStudy (anonymous):

Yes, science, math,physics, or engineering and become a officer thru ROTC, OTS or USAFA

OpenStudy (anonymous):

Oh, right. How old are you?

OpenStudy (anonymous):

13

OpenStudy (anonymous):

Oh, ok, you've got plenty of time to practise maths, then. Just keep asking why.